OpenStudy (mgcanez):

Which expression is equivalent to (7 to the 3rd power) to the -2nd power A.1/7x7x7x7x7x7 B.7 C.1/6x6x6x6x6x6x6 D.1/7

OpenStudy (mhchen):

\[7^{3^{-2}} = \frac{1}{7^{3^{2}}} = \frac{1}{7^{6}} = \frac{1}{7*7*7*7*7*7}\]

OpenStudy (mgcanez):

So it would be A?

OpenStudy (steve816):

Hey there, \[\huge 7^{3^{-2}}=\frac{ 1 }{ 7^6 }\]

OpenStudy (steve816):

That is right! A

OpenStudy (mgcanez):

thank you!

OpenStudy (steve816):

No problem!